OpenStudy (anonymous):

How to show that the left hand side of the below equation equals the right hand side ? \(\dfrac{cos A}{1+sin A} = tan(45+\dfrac{A}{2}) \)

OpenStudy (anonymous):

Any idea ?

OpenStudy (experimentx):

half angle

OpenStudy (badhi):

\[\begin{align*} \frac{\cos A}{1+\sin A}&=\frac{\sin(90+A)}{1-\cos(90+A)}\\ &=\frac{2\sin\left(45+\frac A2\right)\cos\left(90+\frac A2\right)}{2 \sin^2 \left(45+\frac A2\right)}\\&=\cot\left(45+\frac A2\right) \end{align*}\] So your identity is wrong. For example substitute A=90, then, LHS = 0/2 = 0 RHS = tan(45+45) =infinity => not equal.

OpenStudy (anonymous):

Should I take half angles for cos A and sin A too ?

OpenStudy (experimentx):

seems that the identity is wrong.

OpenStudy (anonymous):

Hmm the identity is definitely wrong.